OpenStudy (anonymous):

Please help!!!! Rewrite y = x2 + 14x + 29 in general form. Rewrite y = 3x2 - 24x + 10 in general form.P

OpenStudy (anonymous):

im so confused. i think the form is y=a(x-h)^2+k though

OpenStudy (anonymous):

@dpaInc?? do you know how to do this?

OpenStudy (anonymous):

im so confused. i think the form is y=a(x-h)^2+k though

OpenStudy (anonymous):

general form is when 1 side of the equal sign is a zero.... standard form: \(\large y = x^2 + 14x + 29\) general form: \(\large 0 = x^2 + 14x-y + 29\) vertex form: \(\large y=a(x-h)^2+k \) where (h, k) is the vertex